How would I find the points at which the parabola formed y=x^2 and the line formed by y=2x+3 intersect
Answer by Alan3354(69443)

How would I find the points at which the parabola formed y=x^2 and the line formed by y=2x+3 intersect
---------------
Since both right sides equal y, they equal each other.
x^2 = 2x + 3
Solve that, get the 2 points of intersection.